OpenStudy (vshiroky):

vshiroky Medals 0 Its been a while since I've done this stuff haha... the next one we had was 4x^2/x+4 when x is -2 We got 8

OpenStudy (anonymous):

Well once you plug x in

OpenStudy (anonymous):

\[4(-2)^2/-2\]

OpenStudy (anonymous):

sorry, \[4(-2)^2/(-2)+4\]

OpenStudy (anonymous):

So you would get \[4\times4/2\]

OpenStudy (vshiroky):

16/2 = 8

OpenStudy (anonymous):

omg am i stupid lol you are right it is 8

OpenStudy (vshiroky):

LOL thank you!

OpenStudy (anonymous):

Its 12 over here, that might be why lmao

OpenStudy (astrophysics):

\[\frac{ 4x^2 }{ x+4 } = \frac{ 4(-2)^2 }{ (-2)+4 } = \frac{ 4(-2)(-2) }{ -2 } = 4(-2) = -8\]

OpenStudy (anonymous):

Astrophysics -2 + 4 doesn't equal -2

OpenStudy (vshiroky):

Haha thank you!!

OpenStudy (astrophysics):

I'm dumb to

OpenStudy (astrophysics):

*high five*

OpenStudy (anonymous):

lol we all have our moments *high five*

OpenStudy (astrophysics):

\[\frac{ 4x^2 }{ x+4 } = \frac{ 4(-2)^2 }{ (-2)+4 } = \frac{ 4(-2)(-2) }{ 2 } = \frac{ 16 }{ 2 } = 8\] there you go!

OpenStudy (usukidoll):

for latex use \frac{}{} @whatdoesthismean ex. \[\frac{1}{2} \]

OpenStudy (anonymous):

lol (and yea i noticed to late @UsukiDoll XD)